marie is sailing from an island to a harbor, a distance of 100km. A strong current takes her 13.7 degrees off course for the first 50 km. How far is she now from her destination?
Answer by mananth(16949)

distance = 100 km
the angle = 13.7
..
Let her be away from the by x km
tan 13.7 = x / 100
x= 100 * Tan 13.7
x= 100*2.14
x=214 km